OpenStudy (blake57roger):

A cylinder has a base with a radius of 7 cm and a height of 16 cm. What is the volume of the cylinder? (Use π = 3.14.)

OpenStudy (gorv):

area = pi *r*r * h =pi*7*7*16

OpenStudy (anonymous):

To find the volume you would have to multiply 7*16*3.14